One of the reactions that occurs in a blast furnace, in which iron ore is converted to cast iron, is Fe2O3 + 3CO → 2Fe + 3CO2 Suppose that 1.34 × 103 kg of Fe is obtained from a 2.14 × 103 kg sample of Fe2O3.

Assuming that the reaction goes to completion, what is the percent purity of Fe2O3 in the original sample?
Answer #1

The given balanced reaction is

                               Fe2O3    +             3CO                    →    2Fe            +                3CO2                                  stoichiometric moles 1.0 moles                       3.0 moles                    2.0 moles                      3.0 mole

stoichiometric weights 160 gms                       84 gms                         112 gms                         132 gms

from the balanced equation,

112 grams of Fe is obtained from 160 grams of Fe2O3

1.34*10^3*10^3 grams of Fe is obtained from 1.34*10^6*160/112 = 1914285.71 grams of Fe2O3

1.34 × 10^3 kg of Fe is obtained from a 1.914*10^3 kg sample of Fe2O3.

But the given weight of Fe2O3 is 2.14 × 10^3 kg

the percent purity of Fe2O3 in the original sample = (1.914*10^3*100)/(2.14 × 10^3) = 89.44%